### v2

@@ 2.1 Human Preference Optimization
Another line of research explored direct finetuning on preferred images Dong et al. (2023); Dai et al. (2024) or adopted reinforcement learning from human feedback (RLHF) Black et al. (2024); Fan et al. (2023); Yang et al. (2024); Wallace et al. (2024). Specifically, DDPO Black et al. (2024) formulated denoising as a multi-step decision-making problem. DPOK Fan et al. (2023) incorporated KL regularization as an implicit reward to stabilize the finetuning process. D3PO Yang et al. (2024) and Diffusion-DPO Wallace et al. (2024) extended the concept of DPO Rafailov et al. (2023) to diffusion models by optimizing the respective policies. Different from these methods, CHATS employs two distinct models to separately capture preferred and dispreferred information.
@@ 5.1 Experimental Settings
Models and datasets. We utilize two model families: diffusion models and flow matching models. For diffusion models, we employ Stable Diffusion 1.5 (SD1.5) Rombach et al. (2022) and Stable Diffusion XL-1.0 (SDXL) Podell et al. (2024). For flow matching models, we deploy In-house T2I, a text-to-image generation model optimized for photorealistic images in e-commerce scenarios, based on the DiT architecture Peebles & Xie (2023). We conduct experiments primarily on two preference optimization datasets, Pick-a-Pic v2 (PaP v2) Kirstain et al. (2023a) and OpenImagePreferences (OIP) Data is Better Together (2024) (see Supp. B.1 for more details).
Evaluation. We utilize publicly available benchmark prompts from GenEval Ghosh et al. (2023), DPG-Bench Hu et al. (2024), and HPS v2 Wu et al. (2023). Detailed information about these benchmarks is provided in Supp. B.3. We employ multiple evaluation metrics, including HPS v2 Wu et al. (2023), ImageReward Xu et al. (2024), and PickScore Kirstain et al. (2023b), which are trained on extensive preference datasets, offering robust insights about authentic human preferences. During sampling, by default we keep $s$ and $\alpha$ as 5 and 0.5, respectively.
@@ 5.2 Main Results
In table 1 we report the HPS v2 results over diverse prompt groups as this aesthetic metric is better aligned with human judgment compared to ImageReward and PickScore metrics Wu et al. (2023) (see Supp. C for full results). CHATS consistently surpasses two baselines in all cases, underlining its robustness in handling various visual styles. Moreover, CHATS effectively narrows the aesthetic gaps among different models. For example, the largest discrepancy in average scores across the six groups occurs between SD1.5 and SDXL. Under the Standard setting, this gap is 8.83, but it decreases to 4.41 when adopting CHATS.
@@ 5.3 Identifying Key Components of CHATS
Robustness of $\alpha$. The hyperparameter $\alpha$ in CHATS controls the balance between conditional and unconditional textual embeddings and regulates the contribution of the dispreferred distribution during sampling. We analyze its impact by varying $\alpha$. As shown in Fig. 3 (see Supp. D for more results), both SDXL and In-house T2I models exhibit consistent trends: as $\alpha$ increases from 0 to approximately 0.5, the HPS v2 results improve significantly. Beyond this range, the performance gradually saturates. Notably, compared to the naive CFG case ($\alpha=0$), the proposed sampling strategy leveraging two models consistently enhances aesthetic scores across different model families. These results indicate that $\alpha$ is robust to variations in its value, with the default setting of $0.5$ performing reliably well.
@@ A.4 Theoretical Foundations of CHATS Sampler
Appendix B Full Experimental Settings
B.1 Models and Finetuning Datasets
@@ B.4 Evaluation Metrics
Appendix C Full Aesthetic Results
@@
Appendix D More Ablation on $\alpha$
@@
Appendix E More Qualitative Results

### v3

@@ 2.1 Human Preference Optimization
Another line of research explored direct finetuning on preferred images Dong et al. (2023); Dai et al. (2024) or adopted reinforcement learning from human feedback (RLHF) Black et al. (2024); Fan et al. (2023); Yang et al. (2024); Wallace et al. (2024). Specifically, DDPO Black et al. (2024) formulated denoising as a multi-step decision-making problem. DPOK Fan et al. (2023) incorporated KL regularization as an implicit reward to stabilize the finetuning process. D3PO Yang et al. (2024) and Diffusion-DPO Wallace et al. (2024) extended the concept of DPO Rafailov et al. (2023) to diffusion models by optimizing the respective policies. CHATS departs from these approaches by jointly modeling both preferred and dispreferred distributions with two coupled networks.
Some more recent methods include NPO Zhang et al. (2024), which employs an unlearning based framework to incorporate negative preference signals, and Diffusion-NPO Wang et al. (2025), which trains a negative preference model to substitute the unconditional branch at sampling rather than jointly modeling both distributions. See Section B for further details.
@@ 5.1 Experimental Settings
Models and datasets. We utilize two model families: diffusion models and flow matching models. For diffusion models, we employ Stable Diffusion 1.5 (SD1.5) Rombach et al. (2022) and Stable Diffusion XL-1.0 (SDXL) Podell et al. (2024). For flow matching models, we deploy In-house T2I, a text-to-image generation model optimized for photorealistic images in e-commerce scenarios, based on the DiT architecture Peebles & Xie (2023). We conduct experiments primarily on two preference optimization datasets, Pick-a-Pic v2 (PaP v2) Kirstain et al. (2023a) and OpenImagePreferences (OIP) Data is Better Together (2024) (see Supp. C.1 for more details).
Evaluation. We utilize publicly available benchmark prompts from GenEval Ghosh et al. (2023), DPG-Bench Hu et al. (2024), and HPS v2 Wu et al. (2023). Detailed information about these benchmarks is provided in Supp. C.3. We employ multiple evaluation metrics, including HPS v2 Wu et al. (2023), ImageReward Xu et al. (2024), and PickScore Kirstain et al. (2023b), which are trained on extensive preference datasets, offering robust insights about authentic human preferences. During sampling, by default we keep $s$ and $\alpha$ as 5 and 0.5, respectively.
@@ 5.2 Main Results
In table 1 we report the HPS v2 results over diverse prompt groups as this aesthetic metric is better aligned with human judgment compared to ImageReward and PickScore metrics Wu et al. (2023) (see Supp. D for full results). CHATS consistently surpasses two baselines in all cases, underlining its robustness in handling various visual styles. Moreover, CHATS effectively narrows the aesthetic gaps among different models. For example, the largest discrepancy in average scores across the six groups occurs between SD1.5 and SDXL. Under the Standard setting, this gap is 8.83, but it decreases to 4.41 when adopting CHATS.
@@ 5.3 Identifying Key Components of CHATS
Robustness of $\alpha$. The hyperparameter $\alpha$ in CHATS controls the balance between conditional and unconditional textual embeddings and regulates the contribution of the dispreferred distribution during sampling. We analyze its impact by varying $\alpha$. As shown in Fig. 3 (see Supp. E for more results), both SDXL and In-house T2I models exhibit consistent trends: as $\alpha$ increases from 0 to approximately 0.5, the HPS v2 results improve significantly. Beyond this range, the performance gradually saturates. Notably, compared to the naive CFG case ($\alpha=0$), the proposed sampling strategy leveraging two models consistently enhances aesthetic scores across different model families. These results indicate that $\alpha$ is robust to variations in its value, with the default setting of $0.5$ performing reliably well.
@@ A.4 Theoretical Foundations of CHATS Sampler
Appendix B More Related Works
We note that a recent method Diffusion-NPO Wang et al. (2025) also employs two models for preference optimization. However, CHATS diverges from this approach in several important respects. First, CHATS is motivated by the observation that no existing approach integrates preference optimization and test time sampling within a unified framework to exploit both positive and negative signals, whereas Diffusion-NPO seeks to address limitations of diffusion models by training a dedicated negative preference model to handle unconditional outputs. Second, Diffusion-NPO trains its models sequentially, first fitting a standard DPO-optimized model and then training a negative preference model on inverted preference data. In contrast, CHATS jointly optimizes both preferred and dispreferred distributions under a single objective, eliminating sequential steps and yielding significantly higher training efficiency. Moreover, CHATS demonstrates superior data efficiency by requiring only approximately 7K preference pairs. Finally, at inference CHATS employs a three term guidance scheme grounded in rigorous theoretical analysis (see Sec. A.4) that combines outputs from both models to guide generation toward preferred modes and repel undesirable ones. Diffusion-NPO relies instead on a two term guidance scheme that substitutes the unconditional predictor with the negative preference model. By integrating preference signals consistently during both training and inference, CHATS achieves more coherent alignment and markedly improved image quality.
Appendix C Full Experimental Settings
C.1 Models and Finetuning Datasets
@@ C.4 Evaluation Metrics
Appendix D Full Aesthetic Results
@@
Appendix E More Ablation on $\alpha$
@@
Appendix F More Qualitative Results